OpenStudy (shaw74):

Find the volume of a cone with a diameter of 5 inches and a height that is 3 times the radius. Use 3.14 for pi, and round your answer to the nearest hundredth. 49.06 in3 78.51in3 147.19 in3 392.50 in3

OpenStudy (anonymous):

Use the formula V = 1/3(πr²)h

OpenStudy (shaw74):

ok

OpenStudy (shaw74):

idk how to plug in for that

OpenStudy (shaw74):

radius is 2.5 but how to plug into formula

OpenStudy (shaw74):

height =7.5

OpenStudy (anonymous):

\[\frac{ 1 }{ 3}\pi * \frac{ 5}{ 2} * \frac{ 5 }{ 2}* \frac{ 15 }{ 2}\]= answer

OpenStudy (shaw74):

49.0625

OpenStudy (anonymous):

Yup, now round to the nearest hundredth. :)